OpenStudy (anonymous):

Evaluate the derivative below and state a corresponding integration formula: d/dx [cosx + xsinx]

OpenStudy (anonymous):

= -sin(x) + sin(x) +xcos(x) = xcos(x) the xsin(x) portion uses the product rule d/dx(f(x)*g(x) = f'(x)g(x) +g'(x)f(x)

OpenStudy (anonymous):

\[d/dx(cosx +xsinx)=\int\limits_{}^{}(xcosx) dx\]